Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day02


Crew 100A Log Book for February 13, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: 12:30 pm / 13:30 pm

Brief Narrative of Field Mission Results: We conducted EVAs in two teams to get acquainted with GPS usage, Space Suit, ATVs and Radio Communications. The team that used the ATVs (2 crew members) tested the radio range while the other team (3 crew members) proceeded walking to sites nearby to use the GPS and take some orientation photos. This EVA served mainly to learn about the different equipment and procedures to do while at an EVA. It was a success that resulted in a lot of lessons learned.

EVA Data/Interpretations: n/a

Engineering/Hab Maintenance: Problems with Spirit already mentioned by previous crew.  Need some light bulbs to replace the ones in lower deck.

Webcam feed was fixed for all  6 cameras. West camera in upper deck was realigned to comply with allignment requirements. Cars were moved to the South side of the Hab.

Plans for Tomorrow: Continue with experiments. Try drone at the exterior of the Hab. Conduct EVA to test stress level in the afternoon.

Inventory: none

Support Requested: n/a

Miscellaneous: n/a


Commander's Report

Ivo Ferreira Reporting

After a long day and a good first night of sleep at the MDRS we started our day by setting up three of our experiments. A lot of questionnaires were filled regarding the quality of our sleep, stress level and even mood. Kévin and Jeffrey took their first temperature pills that are supposed to stay inside them until tonight for our sleep study. Let's hope our meals don't make their stomach turn...  Ayako measured our heart pressure and made sure everyone was relaxed before starting up our day. Then we had to take some saliva samples and all this before breakfast!

After breakfast we did our first EVA. The goal was to get acquainted with the Space Suit, ATVs and Radio Communications. We managed to get everyone involved with 5 members in the field and Jeffrey back in the Habitat. Audrey and Romain departed on a small ATV trip with the goal of finding the Radio range. Ayako, Kévin and I did a walking trail to get used to the GPS and document some photos. A lot of lessons learned for the upcoming days...

The afternoon was reserved to set up some other experiments, including a teleoperation with a Rover (Max5J) developed by NASA for environments such as this one. Our expert Jeffrey helped in setting everything up and we managed to commission all the equipment. After some hurdles we managed to get everything up and running.

The scheduled is defined for the following 2 weeks and it looks pretty full.

Everyone is anxious to get some good data from our experience.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.85

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.63

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 12 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): almost Full

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 12

· Potable Water Meter Reading: 420019

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: N/a

· Condition of plants in Tank 2: Ok

· Condition of plants in Tank 3: Ok

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) Yes

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 74

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 64


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All is well


General Engineering

1. Notes/Comments:


EVA 1 Report

Ivo Ferreira Reporting

Crew Members: Ivo Ferreira, Audrey Lan San, Romain Benchenafi, Ayako Ono, Kévin Théatre

Site Location:  Foot: 518567,4250453; ATV: 518882, 4251272

Transit Mode: ATVs, foot

Objectives: Get acquainted with ATV, GPS, Radio Communications and Spacesuit

Results: Managed to comply with the objective with a lot of lessons learned.

Lessons Learned:  Don't use thigh glasses when doing EVA. They get foggy quite easily. Communication should be done with a protocol starting with the name of the person to address, followed by the person that wants to communicate and then the message. The radio coverage is quite large if there are no geographical hurdles in between.


Journalist Report

Romain Benchenafi Reporting

Our second day in the MDRS was for our crew the occasion to make some tests for our future experiments. First we tried an EVA in which all the crew has been involved. It was the occasion to drive ATVs and to discover the surroundings of the habitat. Secondly we tried the Rover, which is working well. That promises good things to come!